2. Find all 6 solutions (real and imaginary) to the equation. x6 = 64
using synthetic division please

using complex zeros
Answer #1

Let x6 = 64 = 64 (140) » XE = 64( cas ofisino) x= 2 ( casoti singo Using Demoiures theorem =>x=, al cos (873) +isin ( 2 ) whput x=3 >*3=2(-14 0i)=-2 put A=4 | 23(-5) - - - - - Ft x= r >x=0/- i) = 1-3 i X, X, , Tg, X3, Xy and as are 6 solutions ano